1/4x-3/8(3x-1/2)=-x-1/8
Answer by faceoff57(108)

1/4x-3/8(3x-1/2)=-x-1/8
1/4x-9/8x+3/16=-x-1/8
16(1/4x-9/8x+3/16=-x-1/8)
4x-18x+3=-16x-2
-14x+3=-16x-2
2x=-5
x=-5/2